A random variable X has the probability density function given below
2(
1
Expert's answer
2020-12-17T19:02:54-0500

X has the probability density function given below


"p(x) = \begin{cases}
 2(1-x) &\text{if } 0\le x < 1 \\
 0, &\text{otherwise }
\end{cases}"

Find "P(X\le2)"


Solution


"P(X<2) = 1-P(X\ge 2) = 1-0 = 1"